If f(x)=5x-1, find the following f(a-2)
Found 2 solutions by AnlytcPhil, psbhowmick:
Answer by AnlytcPhil(1810)

If f(x) = 5x - 1, find the following f(a-2)


f(a number or expression) = right hand side

just means to substitute that number or 
expression for x in the right hand side
and then simplify it:

f(x) = 5x - 1

Erase all the x's and replace them with blanks

f___ = 5_ - 1

Make the blanks long enough to hold (a-2)

f_____ = 5_____ - 1


Put (a-2)'s in the blanks.

f(a-2) = 5(a-2) - 1

Erase the blanks:

f(a-2) = 5(a-2) - 1

Simplify the right side:

f(a-2) = 5a - 10 - 1

f(a-2) = 5a - 11

Edwin


Answer by psbhowmick(878)

f(x) = 5x - 1
To get f(a-2), replace 'x' with (a-2).
Then you get, f(a-2) = 5(a-2) - 1 = 5a - 10 - 1 = 5a - 11